Q.) The temp. Of earth is maintained by a dynamic equilibrium b/w sun and earth.sun & earth can be assumed to be black bodies :A) If the power output of sun would double with changing the temperature,equilibrium temp. Of earth also doubles.B) If the radius of sun doubles without changing it`s power ,it`s surface temperature would decrease by factor of √2 .C) If the radius of Earth devils without any change in Sun its equilibrium temperature wood increase by factor of root 2.D) if there distance between earth and Sun wood decrease by a factor of 2 the equilibrium temperature of earth would increase by factor of root 2.

Grade:12

3 Answers

samm
6 years ago
First calculate the area and then the flux (power). The area is A = 4πR2 A = 4(3.1416)(7 x 108)2 A = 6.16 x 1018 m2 The flux is W = σ T4 W = 5.6704 × 10-8 (5780)4 W = 6.33 x 107 Watts/m2 The total power is P = WA= (5.19 x 107 Watts/m2)(6.16 x 1018 m2 ) P = 3.9 x 1026 Watts
 
samm
6 years ago
We use the distance from the earth to the sun to obtain the flux at the earth. The earth is Re = 1.5 x 108 km from the sun. The area irradiated is Ae = 4πRe 2 Ae = 2.83 x 1023 m2
 
samm
6 years ago
The flux in space above the earth is called the insolation. The insolation is the power coming from the sun divided by the total area at the radius of the earth. We = P/Ae = (3.9 x 1026 Watts)/(2.83 x 1023 m2 ) We = 1.37 x 103 Watts/m2 This is very close to the measured value for radiation in space above the earth.